40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 9 15:31:45 2017, gautam, Update, IMC, IMC wonkiness IMG_7409.JPG
    Reply  Fri Jun 9 17:45:21 2017, Gautam, Kaustubh, Update, IMC, IMC wonkiness 
       Reply  Fri Jun 9 19:18:10 2017, gautam, Update, IMC, IMC wonkiness MC2_UL_glitchy.pngMC2_glitch_fast.png
          Reply  Mon Jun 12 22:23:20 2017, rana, Update, IMC, IMC wonkiness 
          Reply  Tue Jun 13 08:40:32 2017, Steve, Update, IMC, IMC wonkiness happy_MC.pnglast_glitch.png
Message ID: 13057     Entry time: Fri Jun 9 17:45:21 2017     In reply to: 13055     Reply to this: 13058
Author: Gautam, Kaustubh 
Type: Update 
Category: IMC 
Subject: IMC wonkiness 

 

Quote:

Once Steve restores the MC2 Trans cameras, I will hand-align the IMC again and see if the alignment holds for a few hours. If it does, I will reset all offsets for the WFS loops and see if they hold. In particular, the MC2 transmitted spot centering servo has a long time constant so could be something funny there.

 

Summary:

In order to switch on the angular alignment for the IMC mirrors, we needed to center the laser onto the quad-photodiodes at the IMC and the AS Table(WFS1 and WFS2)

I and Gautam went to the IMC table and did the dc centering for the quad-photodiode by varying the beamsplitter angles. After this, we turned the WFS loops off and performed beam centering for the Quad PDs at the AS Table, the WFS1 and WFS2.

Once we had the beam approximately centered for all of the above 3 PDs, we turned on the locking for IMC, and it seems to work just fine. We are waiting for another hour for switching on the angular allignment for the mirrors to make sure the alignment holds with WFS turned off.

ELOG V3.1.3-